I need an in depth step by step explaination on how to solve these problems by elimination. 7p+5q=2/8p-9q=17 and 2x+y=13/4x+2y=23
Answer by ankor@dixie-net.com(22740)

The goal is to mult one or both equations with values that, when the equations are added or subtracted, one of the unknowns, is eliminated.
:
7p + 5q = 2
8p - 9q =17
:
In this problem, we will choose to eliminate q. We can do this if we mult the 1st equation by 9 and the 2nd equation by 5, we then have:
:
63p + 45q = 18
40p - 45q = 85
---------------- add
103p + 0 = 103
It becomes easy to solve for p
103p = 103
Divide both sides by 103 to get 1p
p = 103/103
p = 1
:
Once we know one unknown, in this case it's p=1, we can substitute for p in either equation and find q, we will use the 1st equation:
7(1) + 5q = 2
7 + 5q = 2
Subtract 7 from both sides:
5q = 2 - 7
5q = -5
Divide equation by 5
q = -5/5
q = -1
:
The solutions: p = +1, q = -1
:
Check the solutions in the 2nd equation, substitute for p & q.
8(+1) - 9(-1) = 18
9 - (-9) = 18
Remember a minus a minus is a +
9 + 9 = 18 proves our solutions
:
:
:
This system of equations has no solution, inconsistant is what it's called
We can see this if we try to eliminate y by mult the 1st equation by 2 and subtracting:
2x + y = 13
4x +2y = 23
:
4x + 2y = 26
4x + 2y = 23
-------------- subtract
0x + 0y = 3, we eliminated both variables
0 = 3, means there is no values for x and y that will satisfy both equations
:
How about this? Can you do this now?